OpenStudy (anonymous):

4x^2-12x/5x-7divide x-3/5x-7

OpenStudy (tamtoan):

(a/b)/(c/d) = (a/b).(d/c) <-- use that formula ...also factor the 4x^2 - 12x ...and divide again...you will get the answer

OpenStudy (anonymous):

I am lost can you help me step by step

OpenStudy (tamtoan):

first, (a/b)/(c/d) = (a/b).(d/c) , use this logic and divide the 2 fractions..what do you get ?

OpenStudy (anonymous):

what is a/b)(c/d)=(a/b)(d/c) what do I divide

OpenStudy (tamtoan):

let's say ...you have ...4/5 divide 5/4 ...what do you get ? and what process did you do to get to the answer ?

OpenStudy (anonymous):

1.25

OpenStudy (tamtoan):

no , it's one fraction divide another ...don't usee calculator, do it with fractions dividing. ...

OpenStudy (anonymous):

ok I am confused

OpenStudy (tamtoan):

read this process that i do with those 2 fractions and apply to your problem ... (4/5) / (5/4) = (4/5) . (4/5) <--- reciprocal of second fraction = (4.4)/(5.5) = 16/25

OpenStudy (anonymous):

so it is simplify to 16/25

OpenStudy (tamtoan):

it's the PROCESS of how to get to the answer is what you need to know in order to solve your problem...

OpenStudy (anonymous):

is the next one 12*7

OpenStudy (tamtoan):

where do you get 12x7 from?

OpenStudy (anonymous):

inthe first fraction 12x,-7

OpenStudy (tamtoan):

can you explain clearly again how you get that?